40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1669     Entry time: Wed Jul 13 20:43:03 2016
Author: Antonio, Andrew 
Type: DailyProgress 
Category: North Cavity 
Subject: Cavity locked!!! 

Motivations
We had problems with the error-signal monitor from the TTFSS board. We thought that something 
was wrong with it because the SNR was very small. Before starting the process of nailing down what was
wrong with the electronics I wanted to put some effort to reach a good lock again because I suspected
that the monitor was wrong. We need to check it (later)!

Conclusions
The cavity now it can be stably locked for half an hour.


 

The cavity was initially easy to lock without the EOM. Few crucial things have been done to make the lock:

  • Lower the power entering the cavity from 3mW to 1mW;
  • Inverter on "-" on the FSS box on the table;

With this changes the lock was possible but the control-signal was very noisy;
 

Thanks to Andrew, he added a lambda/4 before the BB EOM to remove additional circular light (after the PBSO).
The quality improved considerably.

I have then twicked the gain Knobs in order to have a better lock. The lock is satisfying (fo now!). Aidan will
provide us of a feedback on the temperature too. Then the lock will stay longer.


Some settings:

  • It cannot be locked with boost ON.;
  • Inverter on FSS box on "-";
  • Common = 600; Fast = 750; Offset = 550;
    The Offset has been adjusted in a way that after the lock the cavity was on resonance (monitoring the DC channel of the PD in reflection);
     
ELOG V3.1.3-